OpenStudy (sneville):

A fossilized leaf contains 15% of its normal amount of carbon 14. How old is the fossil (to the nearest year)? Use 5600 years as the half-life of carbon 14. Solve the problem. A. 35,828 B. 15,299 C. 1311 D. 21,839

OpenStudy (mhchen):

Any ideas?

OpenStudy (sneville):

im stumped

OpenStudy (mhchen):

So for every half-life, its carbon content decreases by 50%. That's (0.5)^x where x is the number of half-lives. Since each half-life is over a period of 5,600 years, then we do (0.5)^(x/5600) So for each 5600 years passed, the content of carbon decreases by 50%. \[Carbon Percentage = 0.5^{\frac{ x }{ 5600 }}\] Do you know how to solve for x?

OpenStudy (wolf1728):

Here's some formulas:

OpenStudy (tkhunny):

Start: 100% 5600 years: 50% 11200 years: 25% 16800 years: 12.5% I think we now have a good idea. No need to worry about some exact manifestation of just the right solution. Do SOMETHING that makes sense and track it down!

OpenStudy (wolf1728):

We want to solve for time so: elapsed time = half-life * log (begng amt/ending amt) / log(2)

OpenStudy (wolf1728):

elapsed time = 5600 * log (100/15) / log(2) elapsed time = 5600 * log (6.6666666667) / log(2) elapsed time = 5600 * .82390874095 / 0.30102999566 elapsed time = 5600 * 2.7369655942 elapsed time = 15,327 AND here's a calculator to check that answer http://www.1728.org/halflife.htm